OpenStudy (anonymous):

∠LMN and ∠NMO are a linear pair of angles. If ∠LMN = 4x + 3 and ∠NMO = 10x – 5, then find the value of x. (1 point) x = 10 x = 12 x = 14 x = 13

OpenStudy (texaschic101):

Linear angles add up to 180 degrees. 4x + 3 + 10x - 5 = 180 -- combine like terms 14x - 2 = 180 -- add 2 to both sides 14x = 182 -- divide by 14 x = 182/14 x = 13

OpenStudy (anonymous):

o iget it now thanks

OpenStudy (texaschic101):

sure thing :)